OpenStudy (bekkah323):

(working with absolute value) solve. 5Ix+4I <= 10

hartnn (hartnn):

if |a| <= b then a<= b or a >=-b so what about 5|x+4| <=10 ??

OpenStudy (johnweldon1993):

So you want to get rid of everything that is not in the absolute value sign 5|x + 4| <= 10 Divide both sides of the equation by 5..this will get rid of the 5 of the left hand side |x + 4| <= 2 Now you can set up your 2 equations *since this is absolute value* \[x + 4 \le 2\] And \[x + 4 \ge -2\] and solve both equations...

OpenStudy (bekkah323):

why did you switch the signs on the second equation

hartnn (hartnn):

lets say for example |x|<= 4 so, x can take values form -4 to +4 (you'll get this if you know what is the effect of absolute sign) so, its like saying -4<= x <= 4 which breaks to x<=4 or x>= -4 (or in other words, you can say that since we multiply by negative, the -4 (2 became -2), we need to switch the sign)

OpenStudy (bekkah323):

ahhh, that makes sense!!!

OpenStudy (johnweldon1993):

Good explanation @hartnn !!

OpenStudy (bekkah323):

so it would be -2 and -6

hartnn (hartnn):

yup

OpenStudy (bekkah323):

-6<= x <= -2

hartnn (hartnn):

yeah

OpenStudy (johnweldon1993):

There we go...I was looking for that lol...yes that is correct!

OpenStudy (bekkah323):

thanks

hartnn (hartnn):

welcome ^_^